# Stochastic Parameter Decomposition

### v1

@@ Abstract
A key step in reverse engineering neural networks is to decompose them into simpler parts that can be studied in relative isolation. Linear parameter decomposition—a framework that has been proposed to resolve several issues with current decomposition methods—decomposes neural network parameters into a sum of sparsely used vectors in parameter space. However, the current main method in this framework, Attribution-based Parameter Decomposition (APD), is impractical on account of its computational cost and sensitivity to hyperparameters. In this work, we introduce Stochastic Parameter Decomposition (SPD), a method that is more scalable and robust to hyperparameters than APD, which we demonstrate by decomposing models that are slightly larger and more complex than was possible to decompose with APD. We also show that SPD avoids other issues, such as shrinkage of the learned parameters, and better identifies ground truth mechanisms in toy models. By bridging causal mediation analysis and network decomposition methods, this demonstration opens up new research possibilities in mechanistic interpretability by removing barriers to scaling linear parameter decomposition methods to larger models. We release a library for running SPD and reproducing our experiments at https://github.com/goodfire-ai/spd.
@@ 2.2 Optimizing for faithfulness
In theory, we could use any arbitrary causal importance function $\Gamma:X\rightarrow[0,1]^{C\times L}$ to predict causal importance values. In practice, we use a set of independent functions $\Gamma^{l}_{c}$, one for each subcomponent. Each function consists of a very small trained MLP $\gamma^{l}_{c}$ (see Appendix A.1 for its architecture). Each MLP takes as input a scalar, the subcomponent’s ‘inner activation’, $h^{l}_{c}(x):=\sum_{j}V^{l}_{c,j}a^{l}_{j}(x)$, where $\vec{a^{l}(x)}$ is the activation vector in the target model that gets multiplied by weight matrix $W^{l}$. The MLP outputs are passed to a hard sigmoid $\sigma_{H}$ to get a prediction of the subcomponent’s causal importance [^3] We use hard sigmoids rather than standard sigmoids because we would like to make it possible for causal importance values to take values of exactly 00 or $1$ or anywhere in between; a standard sigmoid function only tends toward 00 or $1$ as its input tends toward $\infty$- ∞ or $\infty$ respectively. However, note that hard sigmoids have large regions where gradients are 00. We therefore actually use leaky-hard sigmoids to increase training stability. Outputs may therefore lie slightly outside of the $[0,1]$ range (Appendix A.2).:
@@
The causal importance values $g^{l}_{c}(x)$ are used to sample masks that randomly ablate each subcomponent $m^{l}_{c}(x,g^{l}_{c}(x))\sim\mathcal{U}(g^{l}_{c}(x),1)$. We use the reparametrization trick (Kingma and Welling, 2013)(Equation 3) to allow gradients to be backpropagated through the masks $m^{l}_{c}(x,g^{l}_{c}(x))$ to train the causal importance functions $\Gamma^{l}_{c}$. The causal importance functions can therefore learn to produce masks that better predict the ablatability of each subcomponent on a given datapoint. However, if the causal importance functions were trained using only the $\mathcal{L}_{\text{stochastic-recon}}$ loss, they could perform optimally (but pathologically) by outputting a causal importance value of $g^{l}_{c}(x)=1$ for every subcomponent on every input, since there is no incentive to learn to predict the full extent of the ablatability of the subcomponents. An additional loss is required to encourage the causal importance function to always try to ablate as much as possible. We therefore penalize causal importance values for being above 00 using a $\mathcal{L}_{\text{importance-minimality}}$ loss:
@@ 3 Results
Code to reproduce our experiments can be found at https://github.com/goodfire-ai/spd. Training details and hyperparameters can be found in Appendix A.4. Additional figures and training logs can be found in the WandB report here.
@@
Table 1: Mean Max Cosine Similarity (MMCS) and Mean L2 Ratio (ML2R) with their standard deviations (to $3$ decimal places) between learned parameter subcomponents and the target model weights in the subcomponents found by SPD for the embedding matrix $W$ matrix in the $\text{TMS}_{5-2}$ and $\text{TMS}_{40-10}$ models (Section 3.1) and $\text{TMS}_{5-2+\text{ID}}$ and $\text{TMS}_{40-10+\text{ID}}$ models (Section 3.2). These results indicate that the ground truth mechanisms are recovered perfectly and with negligible shrinkage for all models.
@@ 3.1 Toy Model of Superposition
$\text{MMCS}(W,\{\vec{U_{c}}\vec{V_{c}}^{\top}\})=\frac{1}{m_{2}}\sum_{j=1}^{m_{2}}\max_{c}\left(\frac{\vec{U_{:,c}}V_{c,j}\cdot W_{:,j}}{||\vec{U_{:,c}}V_{c,j}||_{2}||W_{:,j}||_{2}}\right)\,,$ (9)
@@
Figure 2: Plots of (left to right) the columns of the input weight matrix $W$ of the $\text{TMS}_{5-2+\text{ID}}$ model (which can be written as $\hat{x}=\text{ReLU}(W^{\top}IWx+b)$, with a weight matrix $W\in\mathbb{R}^{m_{1}\times m_{2}}$ and an identity matrix $I\in\mathbb{R}^{m_{1}\times m_{1}}$); the sum of the parameter subcomponents for that matrix found by SPD; and the individual parameter subcomponents. Although this run of SPD used $20$ subcomponents, only $6$ subcomponents are shown, ordered by the sum of their matrix norms. The first five have learned one direction each, each corresponding to one of the columns of the target model. The final column and the other 14 components (not shown) have a negligible norm because they are superfluous for replicating the behavior of the target model.
Figure 3: Plots of (left to right) the weights in the hidden identity matrix $I$ of the $\text{TMS}_{5-2+\text{ID}}$; the sum of all subcomponents found by SPD for that matrix (including small-norm subcomponents that are not shown); and the largest three individual subcomponents. We see that SPD finds two subcomponents that together sum to the original rank-$2$ identity matrix of the target model, while the other subcomponents have a negligible weight norm.
@@ 3.3 Toy Model of Compressed Computation
Figure 5: The architecture of the Toy Model of Compressed Computation. It uses a $1$-layer residual MLP. Figure adapted from Braun et al. (2025).
@@
[…]
@@
[…]
@@
Figure 6: Toy Model of Compressed Computation: Similarity between target model weights and SPD subcomponents for the first $10$ (out of $100$) input feature dimensions. Top: Neuron contributions measured by Equation 11 for each input feature index $i\in\{0,\dots,9\}$. Bottom: Neuron contributions for the corresponding parameter subcomponents, measured by Equation 12 for each input feature index $i\in\{0,\dots,9\}$. The neurons are numbered from 00 to $49$ based on their raw position in the MLP layer. An extended version of this figure showing all input features and parameter components can be found here.
@@
Figure 8: Causal importance values of each subcomponent (clipped between 00 and $1$) in response to one-hot inputs ($x_{i}=0.75$) for multiple importance loss coefficients $\beta_{3}$. The columns of each matrix are permuted differently, ordered by iteratively choosing the subcomponent (without replacement) with the highest causal importance for each input feature. When the importance loss coefficient $\beta_{3}$ is too low (Left), subcomponents in $W_{\text{in}}$ are not ‘monosemantic’ (i.e. multiple subcomponents have causal importance for the same feature). When the importance loss coefficient $\beta_{3}$ is just right (middle column), each subcomponent in $W_{\text{in}}$ has causal importance for computing a unique input feature. It also identifies the correct number of subcomponents in $W_{\text{out}}$ ($50$). Although it identifies the correct number of components, these subcomponents need not align with any particular basis, and hence look ‘noisy’ because they align with multiple features. But this does not matter, since they always co-activate together and sum to the target model’s identity-matrix parameters. When the importance loss coefficient $\beta_{3}$ is too high (Right), the rank-$50$ $W_{\text{out}}$ component is split into too many subcomponents—approximately one subcomponent for each feature, but where many of the subcomponents have small causal importance values for other features. Also, the causal importance values on the diagonal shrink far below $1.0$, resulting in high $\mathcal{L}_{\text{stochastic-recon}}$ and $\mathcal{L}_{\text{stochastic-recon-layerwise}}$ losses.
@@ 3.4 Toy Models of Cross-Layer Distributed Representations
Figure 10: Toy Model of Distributed Representations (Two Layers): Similarity between target model weights and SPD subcomponents for all $100$ input feature dimensions in a $2$-layer residual MLP. Each point represents one neuron’s contribution to a particular input feature. X-axis: Neuron contributions measured by Equation 11. Y-axis: Neuron contributions for the same neuron on the same input feature in the corresponding parameter subcomponents of $W^{1}_{\text{in}},W^{2}_{\text{in}}$, measured by Equation 12. There is a close match between the X and Y axes for each neuron contribution, indicating that each subcomponent connects its corresponding feature to the MLP neurons with similar weights as the target model. However, there is a systematic skew toward higher values on the Y-axis, indicating that the neuron contributions of the subcomponents tend to be slightly larger. This is in contrast to the one-layer case (Figure 7) and three-layer case (Figure 11). We currently do not understand the source of this discrepancy, but it is possibly an outcome of suboptimal hyperparameters.
@@
SPD finds qualitatively similar results to the $1$-layer Toy Model of Compressed Computation presented in Section 3.3. In the two-layer model, the MLP input matrices $W^{1}_{\text{in}},W^{2}_{\text{in}}$ are decomposed into $100$ components that each compute the functions for one of the input features, using neurons spread over both MLP layers. The MLP output matrices $W^{1}_{\text{out}},W^{2}_{\text{out}}$ are decomposed into a single rank $50$ component. Notably, however, this component is now spread over both layers (Figure 12). This demonstrates that SPD can identify mechanisms that span multiple layers.
@@
In the three-layer model, the MLP input matrices $W^{1}_{\text{in}},W^{2}_{\text{in}},W^{3}_{\text{in}}$ are likewise decomposed into $102$ components all spread over 3 layers, and the MLP output matrices $W^{1}_{\text{out}},W^{2}_{\text{out}},W^{3}_{\text{out}}$ are decomposed into a single rank $51$ component spread over all three layers (Figure 13). In both the two- and three-layer models, we find that the computations occurring in each parameter subcomponent of the $W_{in}$ matrices closely correspond to individual input feature computations in the target model (Figures 10, 11, 14, 15). We note, however, that there is sensitivity to random seeds in these results for these models, possibly due to unfavorable random initialization, which we intend to explore in future work.
@@
Figure 11: Toy Model of Distributed Representations (Three Layers): Similarity between target model weights and SPD subcomponents for all $102$ input feature dimensions in a $3$-layer residual MLP. Each point represents one neuron’s contribution for a particular input feature. X-axis: Neuron contributions measured by Equation 11. Y-axis: Neuron contributions for the same neuron on the same input feature in the corresponding parameter subcomponents of $W^{1}_{\text{in}},W^{2}_{\text{in}},W^{3}_{\text{in}}$, measured by Equation 12. There is a close match between the X and Y axes for each neuron contribution, indicating that each subcomponent connects its corresponding feature to the MLP neurons with similar weights as the target model.
@@
Figure 12: Toy Model of Distributed Representations (Two Layers): Causal importance values of each subcomponent (clipped between 00 and $1$) for the matrices in both MLP layers, $W^{1}_{\text{in}},W^{1}_{\text{out}},W^{2}_{\text{in}},W^{2}_{\text{out}}$ in response to one-hot inputs ($x_{i}=0.75$). Each subcomponent in $W^{1}_{\text{in}},W^{2}_{\text{in}}$ has causal importance for computing a unique input feature. On the other hand, the combined $50$ subcomponents of $W^{1}_{\text{out}},W^{2}_{\text{out}}$ all coactivate for all input features, indicating they are part of a single rank-$50$ identity component. The columns of each matrix are permuted differently, ordered by iteratively choosing the subcomponent (without replacement) with the highest causal importance for each input feature.
@@
Figure 13: Toy Model of Distributed Representations (Three Layers): Causal importance values of each subcomponent (clipped between 00 and $1$) for the matrices in all three MLP layers, $W^{1}_{\text{in}},W^{1}_{\text{out}},W^{2}_{\text{in}},W^{2}_{\text{out}},W^{3}_{\text{in}},W^{3}_{\text{out}}$ in response to one-hot inputs ($x_{i}=0.75$). Each subcomponent in $W^{1}_{\text{in}},W^{2}_{\text{in}},W^{3}_{\text{in}}$ has causal importance for computing a unique input feature. On the other hand, the combined $51$ subcomponents of $W^{1}_{\text{out}},W^{2}_{\text{out}},W^{3}_{\text{out}}$ all coactivate for all input features, indicating they are part of a single rank-$51$ identity component. The columns of each matrix are permuted differently, ordered by iteratively choosing the subcomponent (without replacement) with the highest causal importance for each input feature.

### v2

@@ Abstract
A key step in reverse engineering neural networks is to decompose them into simpler parts that can be studied in relative isolation. Linear parameter decomposition—a framework that has been proposed to resolve several issues with current decomposition methods—decomposes neural network parameters into a sum of sparsely used vectors in parameter space. However, the current main method in this framework, Attribution-based Parameter Decomposition (APD), is impractical on account of its computational cost and sensitivity to hyperparameters. In this work, we introduce Stochastic Parameter Decomposition (SPD), a method that is more scalable and robust to hyperparameters than APD, which we demonstrate by decomposing models that are slightly larger and more complex than was possible to decompose with APD. We also show that SPD avoids other issues, such as shrinkage of the learned parameters, and better identifies ground truth mechanisms in toy models. By bridging causal mediation analysis and network decomposition methods, this demonstration opens up new research possibilities in mechanistic interpretability by removing barriers to scaling linear parameter decomposition methods to larger models. We release a library for running SPD and reproducing our experiments at https://github.com/goodfire-ai/spd/tree/spd-paper.
@@ 2.2 Optimizing for faithfulness
In theory, we could use any arbitrary causal importance function $\Gamma:X\rightarrow[0,1]^{C\times L}$ to predict causal importance values. In practice, we use a set of independent functions $\Gamma^{l}_{c}$, one for each subcomponent. Each function consists of a very small trained MLP $\gamma^{l}_{c}$ (see Appendix A.1 for its architecture). Each MLP takes as input a scalar, the subcomponent’s ‘inner activation’, $h^{l}_{c}(x):=\sum_{j}V^{l}_{c,j}a^{l}_{j}(x)$, where $\vec{a^{l}(x)}$ is the activation vector in the target model that gets multiplied by weight matrix $W^{l}$. The MLP outputs are passed to a hard sigmoid $\sigma_{H}$ to get a prediction of the subcomponent’s causal importance [^3] We use hard sigmoids rather than standard sigmoids because we would like to make it possible for causal importance values to take values of exactly 0 or $1$ or anywhere in between; a standard sigmoid function only tends toward 0 or $1$ as its input tends toward -$\infty$ or $\infty$ respectively. However, note that hard sigmoids have large regions where gradients are 0. We therefore actually use leaky-hard sigmoids to increase training stability. Outputs may therefore lie slightly outside of the $[0,1]$ range (Appendix A.2).:
@@
The causal importance values $g^{l}_{c}(x)$ are used to sample masks that randomly ablate each subcomponent $m^{l}_{c}(x,g^{l}_{c}(x))\sim\mathcal{U}(g^{l}_{c}(x),1)$. We use the reparametrization trick (Kingma and Welling, 2013)(Equation 3) to allow gradients to be backpropagated through the masks $m^{l}_{c}(x,g^{l}_{c}(x))$ to train the causal importance functions $\Gamma^{l}_{c}$. The causal importance functions can therefore learn to produce masks that better predict the ablatability of each subcomponent on a given datapoint. However, if the causal importance functions were trained using only the $\mathcal{L}_{\text{stochastic-recon}}$ loss, they could perform optimally (but pathologically) by outputting a causal importance value of $g^{l}_{c}(x)=1$ for every subcomponent on every input, since there is no incentive to learn to predict the full extent of the ablatability of the subcomponents. An additional loss is required to encourage the causal importance function to always try to ablate as much as possible. We therefore penalize causal importance values for being above 0 using a $\mathcal{L}_{\text{importance-minimality}}$ loss:
@@ 3 Results
Code to reproduce our experiments can be found at https://github.com/goodfire-ai/spd/tree/spd-paper. Training details and hyperparameters can be found in Appendix A.4. Additional figures and training logs can be found in the WandB report here.
@@
Table 1: Mean Max Cosine Similarity (MMCS) and Mean L2 Ratio (ML2R) with their standard deviations (to 33 decimal places) between learned parameter subcomponents and the target model weights in the subcomponents found by SPD for the embedding matrix $W$ matrix in the $\text{TMS}_{5-2}$ and $\text{TMS}_{40-10}$ models (Section 3.1) and $\text{TMS}_{5-2+\text{ID}}$ and $\text{TMS}_{40-10+\text{ID}}$ models (Section 3.2). These results indicate that the ground truth mechanisms are recovered perfectly and with negligible shrinkage for all models.
@@ 3.1 Toy Model of Superposition
[…]
@@
Figure 2: Plots of (left to right) the columns of the input weight matrix $W$ of the $\text{TMS}_{5-2+\text{ID}}$ model (which can be written as $\hat{x}=\text{ReLU}(W^{\top}IWx+b)$, with a weight matrix $W\in\mathbb{R}^{m_{1}\times m_{2}}$ and an identity matrix $I\in\mathbb{R}^{m_{1}\times m_{1}}$); the sum of the parameter subcomponents for that matrix found by SPD; and the individual parameter subcomponents. Although this run of SPD used 2020 subcomponents, only $6$ subcomponents are shown, ordered by the sum of their matrix norms. The first five have learned one direction each, each corresponding to one of the columns of the target model. The final column and the other 14 components (not shown) have a negligible norm because they are superfluous for replicating the behavior of the target model.
Figure 3: Plots of (left to right) the weights in the hidden identity matrix II of the $\text{TMS}_{5-2+\text{ID}}$; the sum of all subcomponents found by SPD for that matrix (including small-norm subcomponents that are not shown); and the largest three individual subcomponents. We see that SPD finds two subcomponents that together sum to the original rank-22 identity matrix of the target model, while the other subcomponents have a negligible weight norm.
@@ 3.3 Toy Model of Compressed Computation
Figure 5: The architecture of the Toy Model of Compressed Computation. It uses a 11-layer residual MLP. Figure adapted from Braun et al. (2025).
@@
$\text{NeuronContribution}_{\text{model}}(i)=({W_{U}}_{[i,:]}W_{\text{out}})\odot(W_{\text{in}}{W_{E}}_{[:,i]})$ (11)
@@
$\text{NeuronContribution}_{\text{subcomponent}}(i)=\max_{c}[({W_{U}}_{[i,:]}U_{\text{out}}V_{\text{out}}^{\top})\odot(U_{\text{in},[:,c]}V_{\text{in},[c,:]}^{\top}{W_{E}}_{[:,i]})]$ (12)
@@
Figure 6: Toy Model of Compressed Computation: Similarity between target model weights and SPD subcomponents for the first $10$ (out of $100$) input feature dimensions. Top: Neuron contributions measured by Equation 11 for each input feature index $i\in\{0,\dots,9\}$. Bottom: Neuron contributions for the corresponding parameter subcomponents, measured by Equation 12 for each input feature index $i\in\{0,\dots,9\}$. The neurons are numbered from 0 to $49$ based on their raw position in the MLP layer. An extended version of this figure showing all input features and parameter components can be found here.
@@
Figure 8: Causal importance values of each subcomponent (clipped between 0 and 11) in response to one-hot inputs ($x_{i}=0.75$) for multiple importance loss coefficients $\beta_{3}$. The columns of each matrix are permuted differently, ordered by iteratively choosing the subcomponent (without replacement) with the highest causal importance for each input feature. When the importance loss coefficient $\beta_{3}$ is too low (Left), subcomponents in $W_{\text{in}}$ are not ‘monosemantic’ (i.e. multiple subcomponents have causal importance for the same feature). When the importance loss coefficient $\beta_{3}$ is just right (middle column), each subcomponent in $W_{\text{in}}$ has causal importance for computing a unique input feature. It also identifies the correct number of subcomponents in $W_{\text{out}}$ ($50$). Although it identifies the correct number of components, these subcomponents need not align with any particular basis, and hence look ‘noisy’ because they align with multiple features. But this does not matter, since they always co-activate together and sum to the target model’s identity-matrix parameters. When the importance loss coefficient $\beta_{3}$ is too high (Right), the rank-$50$ $W_{\text{out}}$ component is split into too many subcomponents—approximately one subcomponent for each feature, but where many of the subcomponents have small causal importance values for other features. Also, the causal importance values on the diagonal shrink far below $1.0$, resulting in high $\mathcal{L}_{\text{stochastic-recon}}$ and $\mathcal{L}_{\text{stochastic-recon-layerwise}}$ losses.
@@ 3.4 Toy Models of Cross-Layer Distributed Representations
Figure 10: Toy Model of Distributed Representations (Two Layers): Similarity between target model weights and SPD subcomponents for all $100$ input feature dimensions in a 22-layer residual MLP. Each point represents one neuron’s contribution to a particular input feature. X-axis: Neuron contributions measured by Equation 11. Y-axis: Neuron contributions for the same neuron on the same input feature in the corresponding parameter subcomponents of $W^{1}_{\text{in}},W^{2}_{\text{in}}$, measured by Equation 12. There is a close match between the X and Y axes for each neuron contribution, indicating that each subcomponent connects its corresponding feature to the MLP neurons with similar weights as the target model. However, there is a systematic skew toward higher values on the Y-axis, indicating that the neuron contributions of the subcomponents tend to be slightly larger. This is in contrast to the one-layer case (Figure 7) and three-layer case (Figure 11). We currently do not understand the source of this discrepancy, but it is possibly an outcome of suboptimal hyperparameters.
@@
SPD finds qualitatively similar results to the $1$-layer Toy Model of Compressed Computation presented in Section 3.3. In the two-layer model, the MLP input matrices $W^{1}_{\text{in}},W^{2}_{\text{in}}$, which each have shape ($\text{d\_mlp}\mathbin{=}25$, $\text{d\_embed}\mathbin{=}1000$), are decomposed into $100$ components that each compute the functions for one of the input features, using neurons spread over both MLP layers. The MLP output matrices $W^{1}_{\text{out}},W^{2}_{\text{out}}$ are decomposed into a single rank $50$ component. Notably, however, this component is now spread over both layers (Figure 12). This demonstrates that SPD can identify mechanisms that span multiple layers.
@@
In the three-layer model, the MLP input matrices $W^{1}_{\text{in}},W^{2}_{\text{in}},W^{3}_{\text{in}}$, which each have shape ($\text{d\_mlp}\mathbin{=}17$, $\text{d\_embed}\mathbin{=}1000$), are likewise decomposed into $102$ components all spread over 3 layers, and the MLP output matrices $W^{1}_{\text{out}},W^{2}_{\text{out}},W^{3}_{\text{out}}$ are decomposed into a single rank $51$ component spread over all three layers (Figure 13). In both the two- and three-layer models, we find that the computations occurring in each parameter subcomponent of the $W_{in}$ matrices closely correspond to individual input feature computations in the target model (Figures 10, 11, 14, 15). We note, however, that there is sensitivity to random seeds in these results for these models, possibly due to unfavorable random initialization, which we intend to explore in future work.
@@
Figure 11: Toy Model of Distributed Representations (Three Layers): Similarity between target model weights and SPD subcomponents for all $102$ input feature dimensions in a 33-layer residual MLP. Each point represents one neuron’s contribution for a particular input feature. X-axis: Neuron contributions measured by Equation 11. Y-axis: Neuron contributions for the same neuron on the same input feature in the corresponding parameter subcomponents of $W^{1}_{\text{in}},W^{2}_{\text{in}},W^{3}_{\text{in}}$, measured by Equation 12. There is a close match between the X and Y axes for each neuron contribution, indicating that each subcomponent connects its corresponding feature to the MLP neurons with similar weights as the target model.
@@
Figure 12: Toy Model of Distributed Representations (Two Layers): Causal importance values of each subcomponent (clipped between 0 and 11) for the matrices in both MLP layers, $W^{1}_{\text{in}},W^{1}_{\text{out}},W^{2}_{\text{in}},W^{2}_{\text{out}}$ in response to one-hot inputs ($x_{i}=0.75$). Each subcomponent in $W^{1}_{\text{in}},W^{2}_{\text{in}}$ has causal importance for computing a unique input feature. On the other hand, the combined $50$ subcomponents of $W^{1}_{\text{out}},W^{2}_{\text{out}}$ all coactivate for all input features, indicating they are part of a single rank-$50$ identity component. The columns of each matrix are permuted differently, ordered by iteratively choosing the subcomponent (without replacement) with the highest causal importance for each input feature.
@@
Figure 13: Toy Model of Distributed Representations (Three Layers): Causal importance values of each subcomponent (clipped between 0 and 11) for the matrices in all three MLP layers, $W^{1}_{\text{in}},W^{1}_{\text{out}},W^{2}_{\text{in}},W^{2}_{\text{out}},W^{3}_{\text{in}},W^{3}_{\text{out}}$ in response to one-hot inputs ($x_{i}=0.75$). Each subcomponent in $W^{1}_{\text{in}},W^{2}_{\text{in}},W^{3}_{\text{in}}$ has causal importance for computing a unique input feature. On the other hand, the combined $51$ subcomponents of $W^{1}_{\text{out}},W^{2}_{\text{out}},W^{3}_{\text{out}}$ all coactivate for all input features, indicating they are part of a single rank-$51$ identity component. The columns of each matrix are permuted differently, ordered by iteratively choosing the subcomponent (without replacement) with the highest causal importance for each input feature.